# InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection

### v2

@@ Abstract
The prevalence of sarcasm in social media, conveyed through text-image combinations, presents significant challenges for sentiment analysis and intention mining. Current multi-modal sarcasm detection methods have been proven to struggle with biases from spurious cues, leading to a superficial understanding of the complex interactions between text and image. To address these issues, we propose InterCLIP-MEP, a robust framework for multi-modal sarcasm detection. InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone, enhancing sample representations by embedding cross-modality information in each encoder. Furthermore, a novel training strategy is designed to adapt InterCLIP for a Memory-Enhanced Predictor (MEP). MEP uses dynamic dual-channel memory to store valuable historical knowledge of test samples and then leverages this memory as a non-parametric classifier to derive the final prediction. By using InterCLIP to encode text-image interactions more effectively and incorporating MEP, InterCLIP-MEP offers a more robust recognition of multi-modal sarcasm. Experiments demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark. Code and data are available at https://github.com/CoderChen01/InterCLIP-MEP.
@@ Introduction
Sarcasm, with its inherent subtlety and complexity, plays a significant role in human communication by often conveying irony, mockery, or hidden intentions (Muecke 1982; Gibbs and O’Brien 1991; Gibbs and Colston 2007). The automatic detection of sarcasm from text has been a significant research focus, aiding tasks such as sentiment analysis and intent mining (Pang, Lee et al. 2008; Tsur, Davidov, and Rappoport 2010; Bouazizi and Ohtsuki 2015; Poria et al. 2016; Amir et al. 2016; Wu et al. 2018; Tay et al. 2018; Baziotis et al. 2018). With the rapid development of social media platforms like Twitter and Reddit, the need for multi-modal sarcasm detection has become increasingly apparent. Users frequently use text-image combinations to convey their messages, making it essential to develop methods to interpret these multi-modal communications. The task of multi-modal sarcasm detection presents significant challenges, requiring the system to accurately capture and comprehend the intricate interplay between textual and visual information to identify sarcasm cues.
A pivotal development in this field was the introduction of the Multi-Modal Sarcasm Detection (MMSD) benchmark by Cai, Cai, and Wan (2019). This benchmark has become a cornerstone for subsequent research, inspiring a range of advanced techniques aimed at improving sarcasm detection performance (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). The above methods have explored various approaches, including the use of Transformer-based models for high-quality representation extraction (Pan et al. 2020), OCR and object detection technologies to enhance image understanding (Qiao et al. 2023), and the incorporation of external emotional knowledge (Liu, Wang, and Li 2022). Graph neural networks have also been utilized to construct dependency and cross-modal graphs for improved sarcasm identification (Liu, Wang, and Li 2022; Liang et al. 2021, 2022).
Figure 1: Comparison of InterCLIP-MEP and other multi-modal sarcasm detection pipelines.
Despite these advancements, the existing methods have shown limitations, often learning biases from spurious cues present in the MMSD benchmark (Qin et al. 2023). To address this, Qin et al. (2023) introduced MMSD2.0, a refined benchmark that eliminates these spurious cues and corrects mislabeled samples. Their work demonstrated the need for more reliable benchmarks and robust detection methods.
Building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection. We present a comparison of existing methods with InterCLIP-MEP in Figure 1. As shown in Figure 1(a) and Figure 1(b), some existing methods use separate uni-modal encoders, such as BERT (Devlin et al. 2019) for text and ViT (Dosovitskiy et al. 2020) for visual modalities, as backbones (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). These encoded samples are fused using various techniques to obtain the final sample representations. We argue that this approach makes it difficult for the model to capture the subtle interactive information between text and images. Although Qin et al. (2023) utilized a multi-modal pre-trained backbone, CLIP (Radford et al. 2021), to encode samples and demonstrated notable results on the MMSD2.0 benchmark, the original CLIP only aligns text and images directly, lacking the nuanced semantic space required for sarcasm detection. Therefore, InterCLIP-MEP introduces a refined variant, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for this specific task. Inspired by the foundational work of Ganz et al. (2024), InterCLIP enhances the capture of interactive information between text and image by embedding representations from one modality into the encoder of the other, leading to a more nuanced understanding of multi-modal sarcasm cues.
@@
After encoding the samples, as depicted in Figure 1(a), many approaches directly identify sarcastic samples through a classification module. This often leads to high prediction entropy and incorrect results, especially for hard cases with greater uncertainty. To the best of our knowledge, few methods consider leveraging historical knowledge of samples to aid in detection. As shown in Figure 1(b), only Wei et al. (2024) constructed a static semantic space that stores representations of all training samples. When identifying sarcastic samples, they query $k$-nearest neighbors in this semantic space and use a voting mechanism to derive the final predictions. However, their static semantic space struggles to handle out-of-distribution issues.
@@
As shown in Figure 1(c), InterCLIP-MEP uses a novel training strategy to adapt InterCLIP for a Memory-Enhanced Predictor (MEP). Based on the sample representations encoded by InterCLIP, InterCLIP-MEP trains a classification module to assign pseudo-labels to samples. Simultaneously, it trains a projection module to map samples into a latent space. In this latent space, same-class samples are pulled closer together, while different-class samples are pushed further apart. Additionally, InterCLIP-MEP fine-tunes the weights of the self-attention modules in InterCLIP’s text and vision encoders using LoRA (Hu et al. 2022), building upon the original CLIP weights. Building on InterCLIP and its training strategy, MEP leverages historical knowledge of test samples to identify sarcasm. It maintains a dynamic, fixed-length dual-channel memory, with one channel storing historical information of non-sarcastic samples and the other storing information of sarcastic samples. MEP uses the learned InterCLIP to extract sample representations, calculate pseudo-labels with the classification module, and compute prediction entropy, while also obtaining projection features through the projection module. Based on the pseudo-labels and prediction entropy, MEP dynamically updates the projected features of samples into the corresponding memory channels, treating these features as the historical knowledge of the test samples. Specifically, MEP consistently updates the memory with projected features of samples that have lower prediction entropy. MEP utilizes this memory as a non-parametric classifier to compute the final prediction for each sample.
Our experiments demonstrate the robustness and innovation of InterCLIP-MEP in accurately identifying sarcasm through sophisticated interaction and memory mechanisms. Our contributions can be summarized as follows:
@@
  • We propose InterCLIP-MEP, a novel framework for reliable multi-modal sarcasm detection, which introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone, improving the encoding of text-image interactions by embedding cross-modality information within each encoder.
  • We design a novel training strategy to adapt InterCLIP for a Memory-Enhanced Predictor (MEP). This strategy enables the integration of dynamic dual-channel memory, which stores valuable historical knowledge of test samples and leverages this memory for more precise and informed predictions.
@@
  • Our extensive experiments on the more reliable benchmark, MMSD2.0, demonstrate that InterCLIP-MEP outperforms existing state-of-the-art methods.
@@
Early work in sarcasm detection primarily focused on textual data (Bouazizi and Ohtsuki 2015; Poria et al. 2016; Amir et al. 2016; Wu et al. 2018; Tay et al. 2018; Baziotis et al. 2018). With the rise of social media, detecting sarcasm from text-image data has become increasingly challenging, leading to the development of several multi-modal approaches (Schifanella et al. 2016; Cai, Cai, and Wan 2019; Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022; Qin et al. 2023; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). Schifanella et al. (2016) were among the first to use multi-modal posts from social platforms to identify sarcasm clues from both text and images. Cai, Cai, and Wan (2019) introduced the MMSD benchmark for multi-modal sarcasm detection, demonstrating its effectiveness with a hierarchical fusion model incorporating image attributes. This benchmark has been widely used in subsequent studies (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022; Qin et al. 2023; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). Further research expanded on MMSD, exploring dependencies and external knowledge, using advanced models, and introducing graph-based methods to enhance performance (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022). However, MMSD has been proven to contain spurious cues that could bias models (Qin et al. 2023). To address this, Qin et al. (2023) introduced MMSD2.0 benchmark, which removed these cues and corrected mislabeled samples. They observed a significant performance drop when re-evaluating state-of-the-art methods on MMSD2.0 and proposed the need for more stable multi-modal sarcasm detection methods. Motivated by the work of Qin et al. (2023), we introduce the InterCLIP-MEP framework for more reliable multi-modal sarcasm detection.
@@
In this paper, we introduce InterCLIP-MEP, a novel framework for multi-modal sarcasm detection that combines interactive cross-modal encoding with memory-enhanced prediction. An overview of InterCLIP-MEP is illustrated in Figure 2. Our approach leverages a variant of CLIP, called Interactive CLIP (InterCLIP), which enhances the original CLIP architecture by embedding representations from one modality into the encoder of the other modality, resulting in more interactive representations. InterCLIP enables a nuanced understanding of the interaction between text and image. To facilitate memory-enhanced prediction, we design a novel training strategy, integrating a classification module and a projection module, with InterCLIP serving as the backbone. These modules are trained to identify sarcastic samples while constructing a latent space where features of the same class are clustered together and those of different classes are distinctly separated. For the prediction phase, InterCLIP-MEP employs a Memory-Enhanced Predictor (MEP) that leverages historical knowledge for final predictions. Instead of relying solely on the classification module to identify sarcasm, MEP compares the features of the current sample with those stored in memory, thereby enhancing the accuracy of sarcasm detection.
@@
We introduce Interactive CLIP (InterCLIP), which can condition text or image encoder to interact with representations from the other modality, enhancing the capture of interactive information crucial for identifying sarcasm.
The input to InterCLIP is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image. Here, for simplicity, we do not consider the case of batch inputs. The text encoder $\mathcal{T}$ extracts the vanilla text representation $\mathbf{F}_{t}$:
@@
Specifically, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in subsequent sections. The structure of the conditioned self-attention layers is illustrated in Figure 3. We denote the input representations to the self-attention layers as $\mathbf{H}_{t/v}$, which are the output representations from the previous layer. The previous layer can either be conditioned or non-conditioned. Due to the dimensional mismatch between the embedded representations $\mathbf{F}_{v/t}$ and the corresponding encoder representation space, we introduce an adapting projection layer $\mathcal{F}_{t/v}$ to project $\mathbf{F}_{v/t}$ into the appropriate representation space.
@@
As shown in Figure 2 (left), to adapt InterCLIP for memory-enhanced prediction, we introduce a novel training strategy. Using InterCLIP as the backbone to obtain fused features of the samples, we introduce a classification module and a projection module.
@@
where $y_{i}$ denotes the label of the $i$-th sample, with sarcastic labeled as 1 and non-sarcastic as 0, and $\hat{y}_{i}$ denotes the prediction for the $i$-th sample. The projection module $\mathcal{F}_{p}$ maps $\tilde{H}^{f}$ into a latent feature space:
@@
where $\textrm{norm}(\cdot)$ denotes L2 normalization, and $d_{f}$ represents the dimension of the projected features. In this space, cosine distance between features of the same class is minimized, while the distance between features of different classes is maximized. We use a label-aware cosine similarity loss to optimize $\mathcal{F}_{p}$:
@@
As depicted in Figure 2 (right), we present a Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance sarcasm detection.
@@
Train 19812 9572 10240
Validation 2410 1042 1368
Test 2409 1037 1372
@@
Table 2: Main results. We use * to indicate that the results are taken from Qin et al. (2023). $\dagger$ indicates the results obtained by running this method on the MMSD2.0 benchmark (Qin et al. 2023).
@@
Param. Value
$r$ 8
top-$n$ 4
$d_{f}$ 1024
$\mathbf{W}$ $W_{k},W_{v},W_{o}$
Epoch 3
Batch size 64
Learning rate 5e-4
LoRA learning rate 1e-4
Table 3: Hyperparameter settings.
Implementation Details
We implement model training and testing using PyTorch Lightning [^1] https://lightning.ai/. We construct InterCLIP leveraging the Transformers library (Wolf et al. 2020), specifically utilizing the clip-vit-base-patch32222https://huggingface.co/openai/clip-vit-base-patch32 model from the HuggingFace Hub [^3] https://huggingface.co/models. We optimize model parameters using AdamW (Loshchilov and Hutter 2017), setting the learning rate to 1e-4 for the LoRA fine-tuning modules and 5e-4 for other trainable modules. We employ a cosine annealing scheduler with warmup to dynamically adjust the learning rate, with warmup steps constituting the first 20% of the total optimization steps and the minimum learning rate set to 1% of the initial rate. For the modules $\mathcal{G}_{t/v}$, $\mathcal{F}_{t/v}$, $\mathcal{F}_{c}$, and $\mathcal{F}_{p}$, we use simple multi-layer perceptrons (MLPs). We perform training with a batch size of 64 for 3 epochs. We run all experiments on a machine with an NVIDIA RTX 4090 GPU.
@@
The main results are shown in Table 2. To validate the effectiveness of our InterCLIP-MEP framework, we conduct experiments using the original CLIP as the backbone instead of InterCLIP, referred to as InterCLIP-MEP w/o Inter. We compare this configuration with three interaction modes of InterCLIP: InterCLIP-MEP w/ V2T, InterCLIP-MEP w/ T2V, and InterCLIP-MEP w/ TW.
For each experiment, we condition only the top four layers of the self-attention modules, with the projection dimension $d_{f}$ set to 1024. We set the rank $r$ of LoRA to 8, fine-tuning the self-attention module weight matrices $\mathbf{W}$, specifically $W_{k}$, $W_{v}$, and $W_{o}$. For the memory size $L$ of the MEP, we select the optimal size from $\mathbf{L}=\{128,256,384,512,640,768,896,1024,1152,1280\}$. Relevant hyperparameters are summarized in Table 3.
Our framework consistently outperforms or matches the performance of state-of-the-art methods, whether using InterCLIP or the original CLIP as the backbone, as shown in Table 2. This demonstrates the effectiveness of our training strategy and MEP. Methods InterCLIP-MEP w/ V2T and InterCLIP-MEP w/ T2V outperform InterCLIP-MEP w/o Inter, indicating that InterCLIP more effectively captures text-image interactions. Additionally, InterCLIP-MEP w/ T2V outperforms InterCLIP-MEP w/ V2T, likely due to the complexity of the visual space, which poses challenges for the projection layer in mapping visual representations into the text encoder space effectively. We observe that InterCLIP-MEP w/ TW performs worse than InterCLIP-MEP w/o Inter, possibly because embedding representations within both encoders increases learning difficulty compared to unidirectional methods. In summary, InterCLIP with T2V interaction, combined with our training strategy and MEP yields the most promising results.
Method Acc. (%) F1 (%)
InterCLIP-MEP w/o Inter
Baseline 86.05 84.81
w/o Proj 85.76 84.43
w/o MEP 85.39 83.99
w/o LoRA 82.44 77.73
InterCLIP-MEP w/ TW
Baseline 85.51 84.26
w/o Proj 85.43 84.05
w/o MEP 85.22 83.79
w/o LoRA 76.42 74.37
InterCLIP-MEP w/ V2T
Baseline 86.26 85.00
w/o Proj 85.68 84.22
w/o MEP 86.26 84.78
w/o LoRA 73.31 72.22
InterCLIP-MEP w/ T2V
Baseline 86.72 85.61
w/o Proj 86.22 84.51
w/o MEP 86.26 84.82
w/o LoRA 75.13 71.79
Table 4: Ablation study results of InterCLIP-MEP. Baseline represents the results of the corresponding method without any ablation.
@@
In this section, we investigate the impact of different hyperparameter settings on model performance. Keeping the other hyperparameters constant as shown in Table 3, we fine-tune all possible weight matrices $\mathbf{W}$ for methods using original CLIP and different interaction modes of Interactive-CLIP as the backbone. We consistently select the optimal memory size from $\mathbf{L}$ for MEP. We calculate the average metrics for different methods and different weight matrices.
$\mathbf{W}$ Mean Acc. (%) Mean F1 (%)
$W_{q}$ 85.14 83.99
$W_{k}$ 85.09 84.00
$W_{v}$ 85.34 84.24
$W_{o}$ 85.39 84.23
$W_{q},W_{k}$ 85.36 84.16
$W_{q},W_{v}$ 85.63 84.40
$W_{q},W_{o}$ 85.73 84.49
$W_{k},W_{o}$ 85.67 84.43
$W_{v},W_{o}$ 85.73 84.54
$W_{k},W_{v}$ 85.70 84.51
$W_{q},W_{k},W_{o}$ 85.92 84.63
$W_{q},W_{v},W_{o}$ 85.87 84.60
$W_{q},W_{k},W_{v}$ 86.05 84.75
$W_{k},W_{v},W_{o}$ 86.14 84.92
$W_{q},W_{k},W_{v},W_{o}$ 85.82 84.55
Table 5: Average results of fine-tuning different weight matrices $\mathbf{W}$ across four baseline methods.
Method Mean Acc. (%) Mean F1 (%)
w/o Inter 85.49 84.32
w/ TW 85.66 84.44
w/ V2T 85.62 84.41
w/ T2V 85.78 84.55
Table 6: Average results of four baseline methods for fine-tuning different weight matrices $\mathbf{W}$.
The results are presented in Table 5 and Table 6. We observe that fine-tuning the weight matrices $W_{k},W_{v},W_{o}$ and using the T2V interaction mode of Interactive-CLIP are the best choices for our InterCLIP-MEP framework.
We further investigate the method using Interactive-CLIP with T2V interaction as the backbone. Keeping the other hyperparameters constant, we condition different top-n layers of the self-attention modules. We also study the impact of different projection dimensions $d_{f}$, different LoRA ranks $r$, and different memory sizes $L$ on the InterCLIP-MEP w/ T2V method. We present all results in Figure 4.

### v3

@@ Abstract
The prevalence of sarcasm in social media, conveyed through text-image combinations, presents significant challenges for sentiment analysis and intention mining. Existing multi-modal sarcasm detection methods have been proven to overestimate performance, as they struggle to effectively capture the intricate sarcastic cues that arise from the interaction between an image and text. To address these issues, we propose InterCLIP-MEP, a novel framework for multi-modal sarcasm detection. Specifically, we introduce an Interactive CLIP (InterCLIP) as the backbone to extract text-image representations, enhancing them by embedding cross-modality information directly within each encoder, thereby improving the representations to capture text-image interactions better. Furthermore, an efficient training strategy is designed to adapt InterCLIP for our proposed Memory-Enhanced Predictor (MEP). MEP uses a dynamic, fixed-length dual-channel memory to store historical knowledge of valuable test samples during inference. It then leverages this memory as a non-parametric classifier to derive the final prediction, offering a more robust recognition of multi-modal sarcasm. Experiments demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark, with an accuracy improvement of 1.08% and an F1 score improvement of 1.51% over the previous best method.
@@ Introduction
Sarcasm, with its inherent subtlety and complexity, plays a significant role in human communication by often conveying irony, mockery, or hidden intentions (Muecke 1982; Gibbs and O’Brien 1991; Gibbs and Colston 2007).
Figure 1: Comparative analysis of multi-modal sarcasm detection pipelines. Panels (a), (b), and (c) depict the pipeline comparisons between our InterCLIP-MEP method and other existing methods. Panel (d) utilizes GradCAM to visually demonstrate the superior ability of the InterCLIP backbone over the traditional CLIP backbone in identifying multi-modal sarcasm cues, which are marked with red boxes.
@@
The automatic detection of sarcasm from text has been a significant research focus, aiding tasks such as sentiment analysis and intent mining (Pang, Lee et al. 2008; Tsur, Davidov, and Rappoport 2010; Bouazizi and Ohtsuki 2015). With the rapid development of social media platforms like Twitter and Reddit, users frequently use text-image combinations to convey their messages. Therefore, the need for multi-modal sarcasm detection has become increasingly apparent, which presents significant challenges, requiring the system to capture and comprehend the intricate interplay between textual and visual information to identify sarcasm cues.
Despite the notable advancements made by existing studies (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022) in this field, Qin et al. (2023) have highlighted that the MMSD benchmark (Cai, Cai, and Wan 2019) employed in these works contains spurious cues. These cues lead to the development of biased models, thereby inflating the perceived effectiveness of these models in detecting sarcasm cues from multi-modal data. By introducing MMSD2.0, a refined benchmark that eliminates spurious cues and corrects mislabeled samples, Qin et al. (2023) demonstrates a substantial performance decline of existing methods when evaluated on MMSD2.0. Although they proposed a promising method that employs a Multi-view CLIP for reliable multi-modal sarcasm detection, it still leaves room for improvement, similar to their previous approaches.
@@
As shown in Figure 1(a) and Figure 1(b), some existing methods (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024) use separate uni-modal encoders, such as BERT (Devlin et al. 2019) for text and ViT (Dosovitskiy et al. 2020) for image, as the backbone to encode samples. This approach makes it difficult for the model to capture the subtle interactive information between text and image. Although Qin et al. (2023) utilized a multi-modal pre-trained backbone, CLIP (Radford et al. 2021), to encode samples and demonstrated notable results on the MMSD2.0 benchmark, the original CLIP only aligns text and images directly, lacking the nuanced semantic space required for sarcasm detection. Furthermore, as depicted in Figure 1(a), some approaches (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Wen, Jia, and Yang 2023; Tian et al. 2023) directly classify sarcastic samples based on their representations. This often leads to high prediction entropy and erroneous outcomes, especially for hard cases with greater uncertainty. To the best of our knowledge, only Wei et al. (2024) considers leveraging historical knowledge of samples to aid in detection. As shown in Figure 1(b), they constructed a static semantic space that stores features of all training samples. When identifying sarcastic samples, they query $k$-nearest neighbors in this semantic space and use a voting mechanism to derive the final predictions. However, relying heavily on a static semantic space may limit its ability to adapt to various test samples.
@@
In this paper, building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection. As shown in Figure 1(c), our InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for detecting multi-modal sarcasm. InterCLIP enhances the capture of interactive information between text and image by embedding representations from one modality into the encoder of the other, leading to a more nuanced understanding of multi-modal sarcasm cues. In Figure 1(d), we utilized GradCAM (Selvaraju et al. 2016) visualization to compare the effectiveness of our InterCLIP backbone against the CLIP backbone in detecting multi-modal sarcasm cues. During inference, InterCLIP-MEP introduces a Memory-Enhanced Predictor (MEP) to build dynamically a non-parametric classifier from valuable historical features to identify multi-modal sarcasm. Specifically, we design an efficient training strategy to adapt InterCLIP for MEP. Based on the sample representations encoded by InterCLIP, we train a classification module to assign pseudo-labels to samples. Simultaneously, we train a projection module to map samples into a latent space. In this latent space, same-class samples are pulled closer together, while different-class samples are pushed further apart. Additionally, we fine-tune the weights of the self-attention modules in InterCLIP’s text and vision encoders using the low-rank adaptation (LoRA) technique (Hu et al. 2022), building upon the original CLIP weights. This approach allows for more efficient adaptation without significantly increasing the computational burden. Building on InterCLIP and the training strategy, MEP leverages historical knowledge of valuable test samples to identify multi-modal sarcasm. It maintains a dynamic, fixed-length dual-channel memory, with one channel storing historical information of non-sarcastic samples and the other storing information of sarcastic samples. MEP then utilizes this memory as a non-parametric classifier to compute the final prediction for each sample, offering more robust and reliable multi-modal sarcasm detection.
Overall, our contributions can be summarized as follows:
@@
  • We propose InterCLIP-MEP, a novel framework for reliable multi-modal sarcasm detection, which introduces Interactive CLIP (InterCLIP) improving the encoding of text-image interactions by embedding cross-modality information within each encoder and Memory-Enhanced Predictor (MEP) offering more robust and reliable multi-modal sarcasm detection.
  • We develop an efficient training strategy that ensures computational efficiency and superior performance, bridging InterCLIP and MEP.
@@
  • Our extensive experiments on the more reliable benchmark, MMSD2.0, demonstrate that InterCLIP-MEP [^1] https://github.com/CoderChen01/InterCLIP-MEP outperforms existing state-of-the-art methods.
@@
Multi-modal sarcasm detection. Early work in sarcasm detection primarily focused on textual data (Bouazizi and Ohtsuki 2015; Amir et al. 2016; Baziotis et al. 2018). With the rise of social media, detecting sarcasm from text-image data has become increasingly challenging, leading to the development of several multi-modal approaches. Schifanella et al. (2016) were among the first to use multi-modal posts from social platforms to identify sarcasm clues from both text and images. Cai, Cai, and Wan (2019) introduced the MMSD benchmark for multi-modal sarcasm detection, demonstrating its effectiveness with a hierarchical fusion model incorporating image attributes. This benchmark has been widely used in subsequent studies (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022; Qin et al. 2023; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). However, MMSD has been proven to contain spurious cues that could bias models (Qin et al. 2023). To address this, Qin et al. (2023) introduced the MMSD2.0 benchmark, which removed these cues and corrected mislabeled samples. They observed a significant performance drop when re-evaluating state-of-the-art methods on MMSD2.0 and proposed the need for more stable multi-modal sarcasm detection methods. In this work, we introduce the InterCLIP-MEP framework towards more reliable multi-modal sarcasm detection.
CLIP adaptation. The Contrastive Language-Image Pretraining (CLIP) model (Radford et al. 2021) excels in vision-language tasks. Adapting CLIP for specific domains has shown substantial improvements, as demonstrated by Li, Shakhnarovich, and Yeh (2022) for phrase localization, Liang et al. (2023) for open-vocabulary semantic segmentation, and Wang et al. (2023) for action recognition. In this work, inspired by Ganz et al. (2024), we conditionally enhance both the text and visual encoders of CLIP, making it more effective in capturing the interplay between text and images to identify sarcasm. Unlike Ganz et al. (2024), who focused solely on embedding text into the visual encoder, we also explore embedding images into the text encoder. Furthermore, their approach is limited to general classification tasks and does not address the complexities of multi-modal sarcasm detection.
Mermory-enhanced prediction. Inspired by cognitive science (Stokes 2015; Baddeley 2000), memory has been introduced to enhance neural networks (Weston, Chopra, and Bordes 2014; Sukhbaatar et al. 2015). Several studies (Wu et al. 2018; Wen, Jia, and Yang 2023) have used memory mechanisms to improve model training, and some (Zhang et al. 2024; Wei et al. 2024) leverage memory to store historical knowledge, enhancing prediction accuracy. In this work, we introduce a memory-enhanced predictor for multi-modal sarcasm detection. In contrast to other methods, our memory dynamically updates during testing, utilizing relevant historical information for improved accuracy and robustness.
@@
An overview of InterCLIP-MEP is illustrated in Figure 2. Initially, we elaborate on the Interactive CLIP (InterCLIP) and its training strategy, followed by an in-depth explanation of the Memory-Enhanced Predictor (MEP).
@@
The input to Interactive CLIP (InterCLIP) is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image. Here, for simplicity, we do not consider the case of batch inputs. The text encoder $\mathcal{T}$ extracts the vanilla text representation $\mathbf{F}_{t}$:
@@
In particular, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in the experiment section. The structure of the conditioned self-attention layers is illustrated in Figure 3. We denote the input representations to the self-attention layers as $\mathbf{H}_{t/v}$, which are the output representations from the previous layer. The previous layer can either be conditioned or non-conditioned. Due to the dimensional mismatch between the embedded representations $\mathbf{F}_{v/t}$ and the corresponding encoder representation space, we introduce an adapting projection layer $\mathcal{F}_{t/v}$ to project $\mathbf{F}_{v/t}$ into the appropriate representation space.
@@
As shown in Figure 2 (left), to adapt InterCLIP for MEP, we introduce an efficient training strategy. Using InterCLIP as the backbone to obtain fused features of the samples, we introduce a classification module and a projection module.
@@
where $y_{i}$ denotes the label of the $i$-th sample, with sarcastic labeled as 1 and non-sarcastic as 0, and $\hat{y}_{i}$ denotes the prediction for the $i$-th sample.
The projection module $\mathcal{F}_{p}$ maps $\tilde{H}^{f}$ into a latent feature space:
@@
where $\textrm{norm}(\cdot)$ denotes L2 normalization, and $d_{f}$ represents the dimension of the projected features. In this space, the cosine distance between features of the same class is minimized, while the distance between features of different classes is maximized. We use a label-aware cosine similarity loss to optimize $\mathcal{F}_{p}$:
@@
As depicted in Figure 2 (right), we present the Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance sarcasm detection.
@@
Train 19,812 9,572 10,240
Validation 2,410 1,042 1,368
Test 2,409 1,037 1,372
@@
Table 2: Main results. We use * to indicate that the results are taken from Qin et al. (2023). $\dagger$ indicates the results obtained by running this method on the MMSD2.0 benchmark (Qin et al. 2023). Underlined values represent the best baseline for comparison. Bold values indicate those that surpass the baseline.
w/o Inter w/ TW w/ V2T w/ T2V
Variant Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%)
Baseline 86.05 84.81 85.51 84.26 86.26 85.00 86.72 85.61
w/o Proj 85.76 84.43 85.43 84.05 85.68 84.22 86.22 84.51
w/o MEP 85.39 83.99 85.22 83.79 86.26 84.78 86.26 84.82
w/o LoRA 82.44 77.73 76.42 74.37 73.31 72.22 75.13 71.79
Table 3: Ablation study of InterCLIP-MEP. Baseline represents the results of the corresponding method without any ablation.
@@
To validate the effectiveness of our InterCLIP-MEP framework, we conduct experiments using the original CLIP as the backbone instead of InterCLIP, referred to as InterCLIP-MEP w/o Inter. We compare this configuration with three interaction modes of InterCLIP: InterCLIP-MEP w/ V2T, InterCLIP-MEP w/ T2V, and InterCLIP-MEP w/ TW.
For each experiment, we condition only the top four layers of the self-attention modules, with the projection dimension $d_{f}$ set to 1024. We set the rank $r$ of LoRA to 8, fine-tuning the self-attention module weight matrices $\mathbf{W}$, specifically $W_{k}$, $W_{v}$, and $W_{o}$. For the memory size $L$ of the MEP, we select the optimal size from $\mathbf{L}=\{128,256,384,512,640,768,896,1024,1152,1280\}$.
The main results are shown in Table 2. Our framework consistently outperforms or matches the performance of state-of-the-art methods, whether using InterCLIP or the original CLIP as the backbone, as shown in Table 2. This demonstrates the effectiveness of our training strategy and MEP. Methods InterCLIP-MEP w/ V2T and InterCLIP-MEP w/ T2V outperform InterCLIP-MEP w/o Inter, indicating that InterCLIP more effectively captures text-image interactions. Additionally, InterCLIP-MEP w/ T2V outperforms InterCLIP-MEP w/ V2T, likely due to the complexity of the visual space, which poses challenges for the projection layer in mapping visual representations into the text encoder space effectively. We observe that InterCLIP-MEP w/ TW performs worse than InterCLIP-MEP w/o Inter, possibly because embedding representations within both encoders increases learning difficulty compared to unidirectional methods. In summary, InterCLIP with T2V interaction, combined with our training strategy and MEP yields the most promising results.
@@
In this section, we further investigate the method using Interactive-CLIP with T2V interaction as the backbone. Keeping the other hyperparameters constant, we condition different top-$n$ layers of the self-attention modules. We also study the impact of different projection dimensions $d_{f}$, different LoRA ranks $r$, and different memory sizes $L$ on the InterCLIP-MEP w/ T2V method. We present all results in Figure 4.